Get Help

Online: Request Help
Phone: (707) 826-HELP (4357)
Email: help@humboldt [dot] edu
In Person: Library 120 • Hours
Reset HSU Password
System Status

Services For

System Status Archive

System Issue start Who is affected What happened
Campus Data Network Mar 13 2015 - 06:00 All campus network users Previously when this was connected it caused a network outage affecting campus network users.
Other Mar 15 2015 - 05:00 All Users Scheduled Monthly Maintenance.
Campus Data Network Mar 11 2015 - 13:30 All Network Users We are currently investigating, check back for updates. UPDATE 1 - Most services have been restored. 2 locations are still down - Bret Hart House and Warren House. TNS is working to restore these services and still investigating the cause of the outage. UPDATE 2 - Bret Hart House and Warren House continue to be down and TNS continues to work to restore these services and investigate the cause of the outage. UPDATE 3 - Services at Bret Hart House and Warren House have been restored.
CMS/Peoplesoft Mar 11 2015 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly schedule maintenance window a fix to the Maxient hold process will be installed.
OBI Mar 8 2015 - 14:00 OBI Consumers Upgrade from 11.1.1.5.5 to 11.1.1.7 will update status when complete. Update is complete.
Other Mar 4 2015 - 15:42 vLab System Administrators are investigating the problem UPDATE: We are still working on the problem. UPDATE2: See Workarounds UPDATE3 See Workarounds
Other Feb 26 2015 - 11:15 Citrix vlab vapp xen Attempting to bring Fischer to aggregated link
Network Folders Feb 24 2015 - 19:00 All Folders Users RedHat has issued a Security Update for Samba (SMB) services, that will require the File Sharing service on Folders to be restarted. This should merely take a second, however, if you are copying data during the service restart, you might get an error message.
Other Feb 24 2015 - 17:00 Users of ere-ssh A critical security patch will be applied.
Other Feb 21 2015 - 01:00 HSE 93 - Brero, HSE 75 - Sculpture, HSE 74 - Ceramics, HSE 55 - Balabanis, HSE 91 - Hagopian, HSE 71 - Little Apartments , HSE 54 - Telonicher TNS is rebooting the switch
Other Feb 20 2015 - 14:40 All Users Power Related Event on Server Farm Switch
Other Feb 20 2015 - 13:32 Users attempting to run DARS Audits. ** DARS Audits are once again available and all PeopleSoft applications appear to be behaving normally. Please contact the Help Desk if you are experiencing any problems with these applications. ** We are currently investigating, check back for updates.
Campus Data Network Feb 20 2015 - 11:30 Campus Network users Preparation work for the upcoming Common Network Initiative (CNI) project.
Campus Data Network Feb 19 2015 - 16:15 Campus Network users CSU CNI ALU Project requires singlemode fiber. We are converting core singlemode fiber links over to multimode fiber where possible to prepare for the refresh.
CMS/Peoplesoft Feb 18 2015 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window MP 15.02 will be installed.
Account Center Feb 16 2015 - 13:50 Account Center/Gamma/OBI Datafile corruption occurred on Gamma requiring a full restore of the database. During this process all data on Gamma will be unavailable as well as the web services normally available through Account Center. OBI login capability is effected by this as well. **Restoration still in progress. While you may be able to access Account Center, we are still working to fully restore all services. Further updates will be provided later this morning.** **Commonly used services have been restored for use although it will be until tomorrow morning before all Account Center jobs are running on schedule. If you’re experiencing a disabled account when attempting to login, please contact the Technology Help Desk for assistance. This systat will be updated tomorrow morning when all services are back on schedule.** **All gamma services should be back to normal. If you are experiencing otherwise, please open a desk desk ticket assigned to programmers so we can look into it.**
Other Feb 16 2015 - 11:43 Natural History Museum [Resolved] PGE reports electrical power issues, Techs resolving by approximately 1:15 pm today
Campus Data Network Feb 16 2015 - 11:23 Users in Telonicher TNS is investigating. Resolved, appears to have been a UPS failure.
Other Feb 16 2015 - 11:43 Natural History Museum PGE reports electrical power issues, Techs resolving by approximately 1:15 pm today
Campus Data Network Feb 15 2015 - 08:00 Wired and Wireless Data Network Users As Part of the Common Network Initiative (CNI)Project, we need to check data link redundancy between core network equipment. This check is a prerequisite to begin the project.
Other Feb 15 2015 - 05:00 All Users Scheduled Monthly Maintenance.
CMS/Peoplesoft Feb 11 2015 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window the Oracle Advance Compression will be installed. This requires extended system downtime 6pm-12am.
Campus Data Network Feb 7 2015 - 11:11 Network users in Buck House (House 097) TNS is investigating. Update 2: TNS is still investigating. Update 3: UPS has lost Power. TNS is working to restore. Update 4: UPS has been replaced. Network has been restored.
Other Feb 7 2015 - 02:02 All Nolij Users ITS Sysadmins are investigating. Errors in logs indicate authentication problems. Underlying Oracle Database issue identified. ITS Database Programmers are investigating. Update 2. The database has been restarted and Nolij is available again
CMS/Peoplesoft Feb 4 2015 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window MP 15.03 (ACA Updates) will be installed.
Campus Data Network Jan 30 2015 - 12:07 Anyone trying to connect to HSU network TNS is investigating.[Issue has been resolved by adjusting local computer settings]
Other Feb 1 2015 - 03:30 All Users RedHat has released a patch for the "Ghost" vulnerability. Postponing the patch until the regular February Maintenance Window on 2-15-2015 would be risky.
Other Jan 29 2015 - 18:30 Nolij, Apex, Dars users. Scheduled maintenance for database.
CMS/Peoplesoft Jan 28 2015 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window, MP 15.00, MP 15.01, and a date page for the Field Trip modification will be installed. We will use an extended maintenance window tonight, expected to last the full six hours.
Wireless Network Jan 21 2015 - 08:45 Any wireless users connecting at the Quad (Update 2) this has been resolved and service restored as of 12:00pm Update as of 10am, this is still down and TNS is still investigating the issue. *** TNS is investigating. Will post again when we have more information.
feedback